A math test is worth 100 points and has 30 problems. Each problem is worth either 3 points or 4 points. How many 4-point problems are there?
Answer by jim_thompson5910(35256)

Hint:

x+y = 30 ---> y = 30-x

3x+4y = 100

3x+4(30-x) = 100

3x+120-4x = 100

solve for x, then use this to find y